Hoover Company purchased two identical inventory items. The item purchased first cost $35.00. The item purchased second cost $38.50. Then Hoover sold one of the inventory items for $60. Based on this information:

a:the amount of ending inventory is $38.50 if Hoover uses the LIFO cost flow method.

b:the amount of gross margin is $23.25 if Hoover uses the weighted average cost flow method.

c:the amount of cost of goods sold is $38.50 if Hoover uses the FIFO cost flow method.

d: the amount of cost of goods sold is $35.00 if Hoover uses the LIFO cost flow method.
Answer #1

As per FIFO ie first in fist out concept the goods sold will be the items purchased first ie cost of goods sold will be =$35

Weighted average cost per unit will be=(35+38.5)/2=$36.75

Hence gross margin=Sales-Cost of goods sold

=(60-36.75)=$23.25

Hence correct answer is B.